# Beyond the Eye: A Relational Model for Early Dementia Detection Using Retinal OCTA Images

### v1

@@ 4.3.2 Regional relationship analysis
The regional relationships graph is a new feature brought by PolarNet+, which can help researchers generate a new understanding of the relations of retinal vascular partitions. We visualize the adjacency matrix generated by the network, using Gephi [5], in the form of a graph of the regional relationship, as shown in Fig. 5, based on the ETDRS grid. Meanwhile, we use a heat map to visualize the adjacency matrix, which aims to provide a different observation perspective, as shown in Fig. 5-(a). From the region-based relations learned by the network in relation to EOAD, we could observe a number of relatively closely related regions by colorizing the weights.
@@
From the Fig. 5-(a), it is evident that partitions DVC-SE and SVC-(NI, SI)111SVC-(NI, SI) denotes SVC-NI and SVC-SI, the following is the same. are closely linked, not only to their internal-layer sectors, but also to the other layer areas: this illustrates the relatively significant cooperating effects of the changes in the corresponding regions in the effects caused by EOAD on the captured retina. Meanwhile, some areas, such as DVC-IE and CC-(II, NI, SI), have relatively weak relations with the others, which indicates that the effect of these regions may be independent, even if their regional importance makes sense.
When we turn the view to the matrix heat map (Fig. 5-(a)), we first found the regions of the DVC have relations with each other, what the regions in other layers did less, and nearly all the DVC regions have strong connections with SVC-(SI, NI) and CC-NE, is less the case with because the DVC is located between the SVC and CC in the retinal tissue structure, having connections with the other layers naturally; and in addition the DVC contains large vessels and capillaries, which provide abundant structural information on the retinal vessels [2], reflecting more EOAD tracks.
Jointly observing the aforementioned three visualizations of relations (Fig. 4-(a) and Fig. 5, we found that the DVC-(SI, NI), SVC-(SI, NI), and CC-NE showed a consistent trend — high importance and strong relationship simultaneously. These are all on the nasal and superior sides, where we found notable increases in amyloid -protein (A) forms and novel intraneuronal A oligomers [21].
@@ 4.4 Model parameters analysis
As shown in Table 3, we conducted a series of experiments to evaluate the performances of the named models by feeding them the same data ().
In our implementation for the EOAD/MCI detection problem, PolarNet+ is constructed by repeating the Res3D block and the multi-view module twice respectively. Consequently, the depth of our implementation is relatively shallow, so its size is tiny (3.92M parameters) when compared to the popular universal models. Although the multi-view property and the relationship-analyzing property, which enhance the high detection performance and interoperability, increased the computational load, modern computing architectures, fortunately, have already covered such scaled computational demands. The advantage of the number of parameters highlights the promise of PolarNet+ deployment on low-cost devices, such as those used for community screening, which could increase the possibility of recognizing EOAD/MCI early.
@@ 5 DISCUSSIONS
To further examine the interpretability of PolarNet+ and to obtain more detailed information about the regional relationships of the OCTA layers, we performed more detailed analyses, such as single-layer analysis (shown in Fig. 5-(b)-(d).) and inter-layer analysis.
When we focus on the SVC-DVC section in Fig. 5-(a), we can see that SVC-(SI, NI) connect to DVC-Tx222DVC-Tx denotes DVC-TI and DVC-TE, the following is the same. and DVC-(Sx, Nx), where SVC-SI performs most actively. Turning to the DVC-CC section in Fig. 5-(a), all DVC regions except DVC-Ix have relations with CC-NE, among which the DVC-(SE, Nx) occupy a prominent position. Although SVC and CC are not directly connected to each other at the tissue structure level, some potential connections do exist: the aforementioned active regions SVC-(SI, NI) show a strong connection to CC-(SE, NE), as shown in the SVC-CC section in Fig. 5-(a). To summarize the inter-layer patterns, we can conclude that the SVC-(SI, NI), CC-NE, and the DVC (except DVC-Ix), play a key role in cooperating effects, and these regions are located in the nasal sides and the superior sides, which is almost in line with the pattern observed in our regional importance analysis.
When we observe the single layer, inner-layer regional relationship patterns are clear. Fig. 5-(b) shows that in the SVC, there is only one strong connection between SVC-SI and SVC-NI, which is even weaker than those in the CC, Fig. 5-(d), where CC-NE is connected as a bridge to the CC-(SE, TE, IE), concentrating on the nasal side. This may suggest that the EOAD influences CC more strongly than the SVC. In the inner-DVC, Fig. 5-(c), connections between regions become strongest and densest, concentrating on the superior sides, with the DVC-SE in the leading position, connecting to DVC-(Nx, Tx, SI).
@@
In this section, we demonstrate the novel interpretability of PolarNet+ through the more detailed visualization of regional relationships. We also acknowledge the limitations of this work: that there is no corresponding clinical research diving into the retinal regional relationships in EOAD. Consequently, this work is future-oriented. Moving from PolarNet to PolarNet+ is a continual exploration of the interpretability of AD/MCI detection based on OCTA images. PolarNet+ inherits its predecessor’s ability to compute regional significance and provides views of regional relationships, with which we can obtain a more open and comprehensive view of AD/MCI analysis based on OCTA images.
@@ 6 CONCLUSION
This paper takes one step in addressing the critical EOAD/MCI detection issue through innovative retinal imaging and deep learning methods. Dementia remains a formidable global health challenge, necessitating early and accurate diagnosis for effective interventions. Our research builds upon the fundamental connection between the brain and the eye, capitalizing on the region-based method and physiological traits to explore retinal biomarkers for EOAD/MCI detection and classification. Retinal OCTA imaging overcomes many defects of established techniques in AD screening, as a non-invasive alternative, enabling the detection of subtle retinal vasculature changes that can serve as valuable AD biomarkers.
In conclusion, the contributions of PolarNet+ are substantial, based on a new mapping paradigm for regional relationships, expanding the understanding of AD. The journey from PolarNet to PolarNet+ is a significant stride in pursuing interpretable and accurate analysis based on OCTA images.

### v2

@@ 4.3.2 Regional relationship analysis
The regional relationships graph is a new feature brought by PolarNet+, which can help researchers generate a new understanding of the relations of retinal vascular partitions. We visualize the adjacency matrix generated by the network, using Gephi [5], in the form of a graph of the regional relationship, as shown in Fig. 6, based on the ETDRS grid. Meanwhile, we use a heat map to visualize the adjacency matrix, which aims to provide a different observation perspective, as shown in Fig. 4. From the region-based relations learned by the network in relation to EOAD, we could observe a number of relatively closely related regions by colorizing the weights.
@@
From the Fig. 4, it is evident that partitions DVC-SE and SVC-(NI, SI)111SVC-(NI, SI) denotes SVC-NI and SVC-SI, the following is the same. are closely linked, not only to their internal-layer sectors, but also to the other layer areas: this illustrates the relatively significant cooperating effects of the changes in the corresponding regions in the effects caused by EOAD on the captured retina. Meanwhile, some areas, such as DVC-IE and CC-(II, NI, SI), have relatively weak relations with the others, which indicates that the effect of these regions may be independent, even if their regional importance makes sense.
Upon examining the matrix heat map (Fig. 4), we observed that the regions within the DVC exhibit stronger inter-regional connectivity compared to the regions in other retinal layers. Nearly all DVC regions demonstrate significant relationships with the regions in the SVC, particularly in the superior and nasal inferior quadrants, as well as with the regions in the CC, particularly in the NE quadrant. This is likely due to the anatomical positioning of the DVC between the SVC and CC within the retinal structure, naturally facilitating interactions with both adjacent layers. Furthermore, the DVC contains larger blood vessels and capillaries, which provide rich structural information regarding retinal vasculature, thereby offering more discriminative patterns for EOAD detection, as suggested by prior studies [2].
A detailed examination of the matrix heat map (Fig. 4) reveals that the regions within the DVC demonstrate a greater degree of inter-regional connectivity in comparison to the regions observed in other retinal layers. The majority of regions in the DVC layer exhibit notable correlations with those in the SVC, particularly in the superior and nasal inferior quadrants, as well as with regions in the CC, particularly in the NE quadrant. This may be attributed to the anatomical positioning of the DVC within the retinal structure, situated between the SVC and the CC. This natural configuration enables interaction with the adjacent retinal layers. Moreover, the DVC contains a greater density of blood vessels and capillaries, which provide detailed structural information regarding the retinal vasculature. This offers a greater number of discriminative patterns for the detection of EOAD, as previously suggested by a clinical study [2].
A joint observation of Fig. 5-(a) and Fig. 6 revealed that the DVC-(SI, NI), SVC-(SI, NI), and CC-(NE) exhibited higher importance weights and a stronger relationships than the other variables. These findings are consistent with those of a previous study [22], which identified notable increases in amyloid -protein (A) forms and novel intraneuronal A oligomers on the nasal and superior sides.
4.4 Effectiveness of polar transformation
To evaluate the effectiveness of the polar transformation, we also performed the classification of different models with and without the polar transformation, and these results are presented in Table 3.
It is evident that the majority of state-of-the-art models have demonstrated enhanced detection performance following the application of the polar transformation of the ETDRS grid, as evidenced by the improvement across all evaluation metrics. In comparison, our PolarNet+ demonstrates a more substantial improvement in performance. While the spatial positional features may undergo alterations following transformation, the key vascular features pertinent to our objective remain largely preserved within the ETDRS grid, and the improvement could be attributed to the transformation’s ability to normalize vessel orientation and spatial relationships.
Furthermore, the vascular alterations resulting from retinal disease are captured by the designed approximate sector convolution (polar transformation), which is intended to emulate the decision-making process frequently employed in clinical practice. This shows that the model focuses on global and local patterns rather than specific spatial relationships, which explains its robustness to such distortions.
The results presented in Table 3 indicate that the classification performance is not affected by the cropping of pixels. This finding aligns with the clinical studies [16, 46, 43], which identified that the pathological changes associated with AD/MCI are predominantly located within the parafoveal region with a specific radius.
Methods | Trans | ACC (meanstd) | AUC (meanstd) | Kappa (meanstd) |
ResNet-34[18] | 0.80060.0146 | 0.80970.0143 | 0.58910.0297 | |
0.83130.0088 | 0.82930.0188 | 0.64910.0238 | ||
EfficientNet-B3[35] | 0.77870.0154 | 0.79450.0088 | 0.54260.0285 | |
0.80390.0194 | 0.79920.0088 | 0.59830.0383 | ||
ConvNeXt-S[27] | 0.78200.0062 | 0.77820.0116 | 0.54850.0165 | |
0.80500.0285 | 0.78840.0169 | 0.59940.0605 | ||
[30] | 0.77540.0142 | 0.79800.0193 | 0.53930.0298 | |
0.79850.0126 | 0.79910.0078 | 0.58440.0285 | ||
VAN-B6[15] | 0.78320.0181 | 0.79490.0187 | 0.55370.0364 | |
0.81490.0124 | 0.81480.0122 | 0.61630.0254 | ||
ViT-Base[12] | 0.62980.0264 | 0.63290.0078 | 0.25020.0412 | |
0.64840.0159 | 0.67840.0218 | 0.28900.0314 | ||
SwinV2-T[26] | 0.66160.0151 | 0.67640.0126 | 0.31190.0271 | |
0.69330.0106 | 0.71180.0192 | 0.37990.0195 | ||
Early fusion[20] | 0.80390.0174 | 0.81820.0096 | 0.59510.0366 | |
0.83680.0179 | 0.83640.0137 | 0.65920.0341 | ||
Middle fusion [48] | 0.81050.0146 | 0.81270.0189 | 0.60430.0287 | |
0.84560.0243 | 0.83580.0133 | 0.68110.0482 | ||
Late fusion[19] | 0.81600.0113 | 0.81750.0136 | 0.62290.0218 | |
0.84770.0179 | 0.83900.0130 | 0.68310.0353 | ||
MCC[47] | 0.77220.0156 | 0.80320.0218 | 0.52590.0339 | |
0.79520.0179 | 0.78520.0083 | 0.57900.0348 | ||
MUCO-Net[38] | 0.83140.0170 | 0.82860.0096 | 0.64900.0296 | |
0.84120.0093 | 0.84620.0136 | 0.66760.0174 | ||
PolarNet[24] | 0.81710.0113 | 0.82890.0095 | 0.62530.0211 | |
0.84890.0187 | 0.85050.0137 | 0.68820.0355 | ||
PolarNet+ | 0.85320.0088 | 0.86050.0055 | 0.69590.0165 | |
0.88390.0122 | 0.88690.0059 | 0.75650.0235 |
@@ 4.5 Model parameters analysis
As shown in Table 4, we conducted a series of experiments to evaluate the performance of the named models by feeding them the same data ().
In our implementation for the EOAD/MCI detection problem, PolarNet+ is constructed by repeating the Res3D block and the multi-view module twice respectively. Consequently, the depth of our implementation is relatively shallow, so its size is tiny (3.92M parameters) when compared to the popular universal models. Although the multi-view property and the relationship-analyzing property, which enhance the high detection performance, increase the computational load, modern computing architectures, fortunately, have already covered such scaled computational demands. The advantage of the number of parameters highlights the promise of PolarNet+ deployment on low-cost devices, such as those used for community screening, which could increase the possibility of recognizing EOAD/MCI early.
@@ 5 DISCUSSIONS
To further examine the potential clinical-acceptability of PolarNet+ and to obtain more detailed information about the regional relationships of the OCTA layers, we performed more detailed analyses, such as single-layer analysis (shown in Fig. 6-(a)-(c).) and inter-layer analysis.
When we focus on the SVC-DVC section in Fig. 4, we can see that SVC-(SI, NI) connect to DVC-Tx222DVC-Tx denotes DVC-TI and DVC-TE, the following is the same. and DVC-(Sx, Nx), where SVC-SI performs most actively. Turning to the DVC-CC section in Fig. 4, all DVC regions except DVC-Ix have relations with CC-NE, among which the DVC-(SE, Nx) occupy a prominent position. Although SVC and CC are not directly connected to each other at the tissue structure level, some potential connections do exist: the aforementioned active regions SVC-(SI, NI) show a strong connection to CC-(SE, NE), as shown in the SVC-CC section in Fig. 4. To summarize the inter-layer patterns, we can conclude that the SVC-(SI, NI), CC-NE, and the DVC (except DVC-Ix), play a key role in cooperating effects, and these regions are located on the nasal sides and the superior sides, which is almost in line with the pattern observed in our regional importance analysis.
When we observe the single layer, inner-layer regional relationship patterns are clear. Fig. 6-(a) shows that in the SVC, there is only one strong connection between SVC-SI and SVC-NI, which is even weaker than those in the CC, Fig. 6-(c), where CC-NE is connected as a bridge to the CC-(SE, TE, IE), concentrating on the nasal side. This may suggest that the EOAD influences CC more strongly than the SVC. In the inner-DVC, Fig. 6-(b), connections between regions become the strongest and densest, concentrating on the superior sides, with the DVC-SE in the leading position, connecting to DVC-(Nx, Tx, SI).
@@
5.1 Limitation of this work
We also acknowledge the limitations of this study. We have conducted additional experiments in which the AD, MCI, and normal datasets were merged and evaluated as a multi-class classification problem. The results were as follows: ACC=0.8858 and AUC=0.9350. However, it was observed that the performance of a three-class classification was superior to that of a binary classification. A detailed examination revealed that the underlying cause was data heterogeneity. The OCTA data collected in two cohorts (EOAD and MCI) were obtained using two different OCTA devices. The model can readily distinguish between EOAD and MCI subjects based on the image style, rather than the pathological features observed in OCTA images. This finding motivates the development of a transfer learning model to normalise images collected from disparate cohorts or clinical centres in future research.
It is also difficult at this stage to validate the link between the explored regional relationships and disease pathogenesis. However, the main focus of this work is to develop an end-to-end detection model and to explore regional interactions in retinal OCTA images in the context of neurodegenerative diseases, which is arguably the first attempt in this field of research. The aim is to move from the analysis of individual features to a more comprehensive relational perspective. It is believed that this approach can help or guide clinicians and biologists to focus on relevant anatomical regions and understand the mechanisms of their association, thus providing a new avenue to deepen the understanding of disease mechanisms.
@@ 6 CONCLUSION
In clinical practice, the region-based analysis technique is frequently employed to study OCTA image biomarkers and to understand the correlation with various eye-related diseases. This is typically achieved by utilising a region-based analysis technique such as the ETDRS grid. To this end, our framework integrates the ETDRS and gradCAM, and maps the model’s decision-making to a region-based representation that aligns with clinical practice. This paper takes one step in addressing the critical EOAD/MCI detection issue through innovative retinal imaging and deep learning methods. PolarNet+ inherits its predecessor’s ability to compute regional significance and provides views of regional relationships, with which we can obtain a more open and comprehensive view of AD/MCI analysis based on OCTA images.
Our method extends beyond the mere averaging of data, incorporating regional relationship modelling to elucidate the manner in which different regions interact and contribute collectively to the decision-making process. This offers a more comprehensive view of the model’s predictions, improving the clinically acceptability using the assessment of regional importance. To the best of our knowledge, this work is the first attempt to generate a regional map of significant importance, specifically designed for EOAD/MCI using OCTA images. This task-specific adaptation not only adds clinical relevance but also provides a new level of insight into regional disease patterns. Although gradCAM is a well-established method for visual explanations, averaging outputs within regions is necessary for tasks that rely on region-level analysis. This approach guarantees consistency with the clinical partitioning strategy and enables the model to concentrate on region-specific disease-relevant patterns.
Dementia remains a formidable global health challenge, necessitating early and accurate diagnosis for effective interventions. Our research builds upon the fundamental connection between the brain and the eye, capitalizing on the region-based method and physiological traits to explore retinal biomarkers for EOAD/MCI detection and classification. Retinal OCTA imaging overcomes many defects of established techniques in AD screening, as a non-invasive alternative, enabling the detection of subtle retinal vasculature changes that can serve as valuable AD biomarkers. In conclusion, the contributions of PolarNet+ are substantial, based on a new mapping paradigm for regional relationships, expanding the understanding of AD. The journey from PolarNet to PolarNet+ represents a substantial advancement in the pursuit of clinical-friendly and precise analysis based on OCTA images.
7 Acknowledgment
This work is supported by the National Key Research and Development Program of China (2024YFB3815100), the National Science Foundation Program of China (62422122, 62272444), the Key Research and Development Program of Zhejiang Province (2020C03036, 2023C04017, 2024C03101), Ningbo 2025 S&T Mega project (2021Z134, 2022Z127, 2022Z134), the Key Project of Ningbo Public Welfare Science and Technology (2023S012).